# A Survey of Long-Document Retrieval in the PLM and LLM Era

### v1

@@ 6.2. Scholarly Long-Document Retrieval
Citation- and concept-aware encoders provide high-recall seeds and stable geometry for long-corpus search: SciBERT supplies scientific-domain pretraining model for passage encodings (Beltagy et al., 2019b); SPECTER injects citation signals via triplet training (query paper, cited positive, uncited/hard-negative) to yield document embeddings aligned with scholarly relatedness—well-suited for document-level initial retrieval refined at passage level (Cohan et al., 2020); SciNCL replaces discrete “cited/not” labels with neighborhood-contrastive sampling over a citation-graph embedding, improving global geometry for paper-to-paper search (Ostendorff et al., 2022).
@@
To control cost on long texts, CORANK extracts zero-shot LLM features (categories/sections/keywords) to rerank large candidate pools, then applies full-text reranking only to the top-$M$ candidates—preserving LDR coverage while reducing token and latency budgets (Tian et al., 2025).
@@
In general, LDR for scholarly papers tames length by (i) abstracting queries to concepts/aspects to steer recall, (ii) leveraging citation graphs for long-range candidate proposals, and (iii) delaying full-text interaction to a narrow shortlist—while enforcing span-level attribution to preserve provenance.
@@ 6.4. Cross-Lingual Long-Text Retrieval
A representative benchmark is XOR-QA(Asai et al., 2020), which decouples the query language from the evidence language (e.g., a Japanese query with only English evidence). Since Wikipedia articles are long and multi-sectioned, models must not only retrieve passages across languages but also aggregate them for provenance-aware answers. Building on this benchmark, McCrolin(Limkonchotiwat et al., 2024) introduces a multi-consistency training framework with a teacher–student setup (frozen mUSE as teacher). It enforces cross-lingual semantic consistency and stable ranking via tailored loss functions, showing strong results on long-document retrieval across different encoders.
@@
Although differing in architecture, context length, and optimization strategy, these approaches converge on the core pain points of cross-lingual long-text retrieval—semantic alignment, scattered evidence, and computational efficiency—while offering complementary solutions for cross-lingual long-context scenarios.
@@ 7. Current Challenges and Future Directions
Despite substantial progress, Long-Document Retrieval remains an open and rapidly evolving research frontier. We summarize the most critical challenges and outline promising directions for future work.
@@ 7.1. Key Challenges
(1) Efficient Scaling. Even with recent advances in long-context PLMs/LLMs (e.g., 32K–100K token windows), the quadratic cost of attention remains prohibitive. Sparse attention and compression techniques mitigate costs but often degrade effectiveness, especially in downstream retrieval and QA tasks where fine-grained signals matter. A fundamental challenge is to design architectures or indexing schemes that scale sub-linearly while maintaining retrieval fidelity. So far, we still lack clear “scaling laws” for retrieval accuracy under extreme context lengths.
(2) Relevance Localization and Multi-Granularity. Long documents mix relevant and irrelevant content. Retrievers must both identify small but critical evidence and synthesize document-level signals. Current block-selection strategies (e.g., KeyB, ICLI) only partially solve this and often suffer from query drift or redundant recall. Discourse-aware and structure-aware retrieval remain underexplored, and how to enable reasoning across blocks without breaking context continuity is still an open question.
(3) Data Scarcity. High-quality document-level labels are expensive. Most datasets provide sparse judgments (often one relevant doc per query), making supervised training difficult. Synthetic data generation and weak supervision help, but current methods mostly operate at sentence-level, lacking document- or paragraph-level supervision. Leveraging cross-task signals (e.g., summarization, QA) or implicit user interaction data may be critical to overcome this bottleneck.
(4) Domain and Language Adaptation. Specialized domains (e.g., legal, biomedical) and multilingual LDR require handling domain-specific structures and vocabularies. Existing work such as MLDR-zh highlights the difficulty of transferring from English-centric datasets. Moreover, structured elements like legal clause numbering, medical tables, or formulae are rarely integrated into retrieval models, leading to poor cross-domain generalization.
(5) Interpretability and User Trust. Users must understand why a 100-page document was retrieved. Current neural/LLM models lack faithful attribution; attention maps or post-hoc highlights often fail to reflect actual reasoning. Silent hallucination is a growing risk in LDR-enhanced LLMs. Providing span-level justifications, designing attribution-aware training objectives, and developing user-centered evaluation protocols remain urgent needs.
@@ 7.2. Future Directions
Unified Retrieval–Reading Models. LLMs blur the boundary between retrieval and QA by directly consuming long contexts. Future systems may jointly optimize retrieval and reasoning, aided by context compression (e.g., LongLLMLingua (Jiang et al., 2024)), selective pruning (LI et al., [n. d.]), and abstraction prompts. However, their cost–benefit tradeoffs under large-scale corpora remain unclear, and the risk of retriever-free paradigms replacing retrievers needs careful investigation.
Advances in Long-Context Architectures. Innovations in efficient attention (RoPE (Su et al., 2024), DuoAttention, external memory (Wu et al., [n. d.])) suggest hybrid models that combine local/global attention and hierarchical representations will be central to scaling beyond 100K tokens. Future research may also explore structured or episodic memory modules, bridging symbolic storage (e.g., graphs, tables) with neural context modeling.
Retrieval-Enhanced LLMs. RAG-style frameworks for long documents remain underdeveloped. Most current work applies retrieval in a plug-and-play manner without query refinement. Iterative retrieval–generation (Shao et al., 2023) points to LLMs as active orchestrators, deciding when to retrieve, summarize, or refine queries. Future systems may require multi-round retrieval loops, adaptive granularity, and tighter integration of retrieval signals into LLM reasoning.
Benchmarking and Evaluation. Current metrics (nDCG, Recall) fail to capture long-context challenges such as hallucination, attribution, and multi-hop reasoning. Moreover, most benchmarks contain documents under 5K–10K words, far from real-world cases such as patents, legal codes, or full books. Future benchmarks should incorporate span-level and document-level double annotations, task-aware metrics (e.g., correctness vs. faithfulness), and user-centric measures such as time-to-answer and cognitive load.
Cross-pollination and Multimodality. Insights from summarization, discourse parsing, and knowledge graphs may inspire new indexing schemes. Beyond text, many long documents are multimodal (figures, tables, videos). A key open challenge is alignment: ensuring tables, figures, and captions are jointly indexed and retrievable with text. Benchmarks like VideoWebArena (Jang et al., [n. d.]) highlight the gap—current multimodal LLMs perform far below human level. Integrating layout modeling (e.g., LayoutLM family) and multimodal reasoning could substantially advance this area.
@@ 7.3. Summary
Long-Document Retrieval lies at the intersection of IR efficiency, NLP understanding, and LLM reasoning. We argue that no single paradigm suffices: future progress will hinge on hybrid solutions that combine efficient indexing, structure-aware modeling, and LLM-based reasoning, while ensuring robustness, interpretability, and domain generalization. Equally important is developing realistic benchmarks, user-centered evaluation, and domain-specific adaptation to ensure that LDR research translates into trustworthy and deployable systems.
@@ 8. Conclusion
Looking ahead, we argue that the most promising progress will come from hybrid architectures—systems that fuse the efficiency of classical IR indexing, the semantic depth of neural encoders, and the reasoning capabilities of LLMs. Equally important is the development of next-generation evaluation frameworks that move beyond coarse-grained metrics to capture attribution, faithfulness, and user-centered utility in long-document settings. As these hybrid systems mature, they promise to transform how we interact with large-scale information—moving beyond simple document retrieval towards a future of deep evidence synthesis and automated knowledge discovery at scale.

### v2

@@ 6.2. Scholarly Long-Document Retrieval
Citation- and concept-aware encoders provide high-recall seeds and stable geometry for long-corpus search: SciBERT supplies scientific-domain pretraining model for passage encodings (Beltagy et al., 2019b); SPECTER injects citation signals via triplet training (query paper, cited positive, uncited/hard-negative) to yield document embeddings aligned with scholarly relatedness, which is well-suited for document-level initial retrieval refined at passage level (Cohan et al., 2020); SciNCL replaces discrete “cited/not” labels with neighborhood-contrastive sampling over a citation-graph embedding, improving global geometry for paper-to-paper search (Ostendorff et al., 2022).
@@
To control cost on long texts, CORANK extracts zero-shot LLM features (categories/sections/keywords) to rerank large candidate pools, then applies full-text reranking only to the top-$M$ candidates, preserving LDR coverage while reducing token and latency budgets (Tian et al., 2025).
@@
In general, LDR for scholarly papers tames length by (i) abstracting queries to concepts/aspects to steer recall, (ii) leveraging citation graphs for long-range candidate proposals, and (iii) delaying full-text interaction to a narrow shortlist while enforcing span-level attribution to preserve provenance.
@@ 6.4. Cross-Lingual Long-Text Retrieval
A representative benchmark is XOR-QA(Asai et al., 2020), which decouples the query language from the evidence language (e.g., a Japanese query with only English evidence). Since Wikipedia articles are long and multi-sectioned, models must not only retrieve passages across languages but also aggregate them for provenance-aware answers. Building on this benchmark, McCrolin(Limkonchotiwat et al., 2024) introduces a multi-consistency training framework with a teacher-student setup (frozen mUSE as teacher). It enforces cross-lingual semantic consistency and stable ranking via tailored loss functions, showing strong results on long-document retrieval across different encoders.
@@
Although differing in architecture, context length, and optimization strategy, these approaches converge on the core pain points of cross-lingual long-text retrieval: semantic alignment, scattered evidence, and computational efficiency, while offering complementary solutions for cross-lingual long-context scenarios.
@@ 7. Current Challenges and Future Directions
Despite substantial progress, long-document retrieval (LDR) remains an open and rapidly evolving research frontier. Below we summarize critical challenges and outline promising research directions.
@@ 7.1. Key Challenges
(1) Efficient Scaling. Even with long-context PLMs/LLMs (e.g., 32K–100K windows), end-to-end processing of full documents remains expensive. Sparse/kernelized attention, chunking, and compression reduce cost but can attenuate fine-grained matching signals that matter for retrieval and QA. A core challenge is to design architectures and indexing schemes that scale sub-linearly in document length while preserving retrieval fidelity. Systematic “scaling laws” for accuracy–cost trade-offs at extreme context lengths are still underexplored.
(2) Relevance Localization and Multi-Granularity. Long documents interleave relevant and irrelevant content. Retrievers must localize small yet critical spans and synthesize document-level evidence. Existing block/segment selection (e.g., KeyB/KeyB2, hierarchical selectors, LLM-guided in-context selection) alleviates cost but can suffer from query drift, redundancy, or broken cross-block reasoning. Discourse- and structure-aware retrieval that respects document organization (sections, tables, figures, citations) remains limited.
(3) Data Scarcity. High-quality document-level supervision is costly, making fully supervised learning difficult. Synthetic/weak supervision helps, but is often sentence-level and lacks document- or paragraph-granularity alignment. Leveraging cross-task signals (summarization, QA, citation linking) or implicit interaction signals is a promising route to richer supervision at document scale.
(4) Domain and Language Adaptation. Specialized domains (legal, biomedical, scholarly) and multilingual LDR introduce domain-specific structures and terminology. Transfer from English-centric datasets remains challenging (e.g., MLDR-zh). Structured artifacts, e.g., clause numbering, formulas, tables, layouts are seldom integrated into retrieval models, limiting cross-domain generalization.
(5) Interpretability and User Trust. Users need to understand why a long document is retrieved. Current neural/LLM models offer weak attribution: attention maps or post-hoc highlights may not reflect actual decision signals, and hallucination risks grow when models summarize long evidence. Span-level rationales, attribution-aware training, and user-centered evaluation protocols are urgently needed.
@@ 7.2. Future Directions
Unified Retrieval–Reading Models. LLMs increasingly blur the retrieval–reasoning boundary by consuming long contexts. Jointly optimizing retrieval and reading, aided by context compression (e.g., LongLLMLingua (Jiang et al., 2024)), selective pruning (LI et al., [n. d.]), and abstraction prompts, is promising. Open questions include when retrieval-free pipelines suffice, and how to characterize their cost–benefit at corpus scale.
Advances in Long-Context Architectures. Progress in efficient attention (sparse/kernelized/chunked), improved positional schemes (e.g., RoPE variants (Su et al., 2024)), and external memory (Wu et al., [n. d.]) suggests hybrid designs that combine local/global attention with hierarchical representations. Structured or episodic memory that bridges symbolic storage (graphs, tables, layouts) and neural context modeling is a natural next step beyond 100K tokens.
Retrieval-Enhanced LLMs. RAG for long documents remains under-engineered. Beyond plug-and-play retrieval, systems should support iterative retrieval-generation cycles (Shao et al., 2023), query refinement, and adaptive granularity (document/section/passage). Tighter integration of retrieval signals into LLM reasoning, e.g., retrieval-conditioned decoding, retrieval-aware training, may improve faithfulness and efficiency.
Benchmarking and Evaluation. Conventional metrics (nDCG, Recall) overlook long-context issues such as attribution, multi-hop reasoning, and user effort. Many benchmarks emphasize short passages or moderate-length articles rather than truly long artifacts (patents, legal codes, books). Future resources should combine span-level and document-level annotations, include task-aware metrics (correctness vs. faithfulness), and report user-centric measures (e.g., time-to-answer, cognitive load) alongside latency and throughput.
Cross-pollination and Multimodality. Insights from summarization, discourse parsing, and knowledge graphs can inspire structure-aware indexing. Many long documents are multimodal; aligning tables, figures, and text for joint indexing and retrieval remains open. Benchmarks such as VideoWebArena (Jang et al., [n. d.]) expose large gaps to human performance; integrating layout modeling (e.g., LayoutLM-style features) and multimodal reasoning is a key avenue.
@@ 7.3. Summary
LDR sits at the intersection of IR efficiency, document understanding, and LLM reasoning. No single paradigm suffices: progress will hinge on hybrid solutions that combine sub-linear indexing, structure-aware modeling, and LLM-based reasoning with robustness, interpretability, and domain generalization. Equally important are realistic long-document benchmarks and user-centered evaluation to ensure LDR systems are trustworthy and deployable.
@@ 8. Conclusion
Looking ahead, we argue that the most promising progress will come from hybrid architectures: systems that fuse the efficiency of classical IR indexing, the semantic depth of neural encoders, and the reasoning capabilities of LLMs. Equally important is the development of next-generation evaluation frameworks that move beyond coarse-grained metrics to capture attribution, faithfulness, and user-centered utility in long-document settings. As these hybrid systems mature, they promise to transform how we interact with large-scale information, moving beyond simple document retrieval towards a future of deep evidence synthesis and automated knowledge discovery at scale.